# Matrix Product State on a Quantum Computer

### v1

@@ Abstract
Solving quantum many-body systems is one of the most significant regimes where quantum computing applies. Currently, as a hardware-friendly computational paradigms, variational algorithms are often used for finding the ground energy of quantum many-body systems. However, running large-scale variational algorithms is challenging, because of the noise as well as the obstacle of barren plateaus. In this work, we propose the quantum version of matrix product state (qMPS), and develop variational quantum algorithms to prepare it in canonical forms, allowing to run the variational MPS method, which is equivalent to the Density Matrix Renormalization Group method, on near term quantum devices. Compared with widely used methods such as variational quantum eigensolver, this method can greatly reduce the number of qubits used in local optimization, and thus mitigate the effects of barren plateaus while obtaining better accuracy. Our method holds promise for distributed quantum computing, offering possibilities for fusion of different computing systems.
@@ I Introduction
Solving quantum many-body systems is a fundamental challenge. Classical computers usually have to deal with the exponential growth in size of the system to mimic the quantum effects. As Feynman raised the concept of quantum simulations Feynman (1982), quantum computing begins to show its great power for specific tasks and stimulated the possibility of solving these systems with advantages over the classical computers. Quantum algorithms were proposed for efficient quantum simulations, and are promising for a wide range of applications such as condensed matter physics Anderson (1984); Sachdev (2023), quantum chemistry Cao et al. (2019); McArdle et al. (2020), high-energy physics Zohar et al. (2015); Martinez et al. (2016), materials Monroe et al. (2021); Maskara et al. (2025), drug discovery Wong et al. (2023); Santagati et al. (2024), or combinatory optimization tasks Moll et al. (2018); Harrigan et al. (2021).
@@
As the quantum advantages have been demonstrated on both superconducting and photonic platforms Arute et al. (2019); Zhu et al. (2022); Madsen et al. (2022); Zhong et al. (2021); Deng et al. (2023); Acharya et al. (2025); Gao et al. (2025), quantum computing has developed rapidly and is moving towards the noisy intermediate scale quantum (NISQ) era Preskill (2018). However, taking the practical physical devices into consideration, running large-scale quantum algorithms still meets severe challenges. For example, algorithms such as the variational quantum eigensolver (VQE) Peruzzo et al. (2014); Kandala et al. (2017), are troubled with ‘barren plateaus’ Liu et al. (2022); Wang et al. (2021), and are more time consuming to converge to an accurate result. To mitigate these challenges, various distributed quantum frameworks have been proposed. For example, deepVQE can handle large-scale systems with weak correlations among subsystems. Other methods such as circuit cutting Peng et al. (2020); Tang and Martonosi (2022) or hybrid tree tensor networks Yuan et al. (2021) reduce the sizes of quantum circuits, but requires global optimizations, which limits the efficiency.
@@
Fortunately, the hybrid tensor network provides a general framework for transplanting the classical tensor network method onto quantum devices. Tensor networks are powerful numerical tools for solving fundamental problems in quantum many-body systems. Recently, tensor network methods are applied in quantum information, and become one of the most efficient method for simulating quantum circuits Markov et al. (2018); Villalonga et al. (2019, 2020); Guo et al. (2019); Liu et al. (2021); Pan et al. (2022). As one of the most widely used tensor network structures, Matrix Product State (MPS) supports important applications of finding the ground energy of a strongly correlated 1-D many-body system Östlund and Rommer (1995); Hastings (2007); Liu et al. (2020).
@@
In this work, we propose the idea of quantum Matrix Product State (qMPS), and develop the variational quantum MPS (vqMPS) method for quantum simulation, which can be viewed as the quantum version of the Density Matrix Renormalization Group (DMRG) method White (1992); Schollwöck (2011) for near-term quantum devices. This method makes it possible to run large-scale quantum simulations with local optimizations on small-scale subsystems, and can reach comparable or even better accuracy over VQE using much fewer qubits. This can greatly enhance the efficiency for both physical demonstration and numerical simulation. Also, the local optimizations can be distributed on different systems, providing a framework for the fusion of different physical platforms, or further quantum-classical hybrid computing.
@@ II The quantum MPS
A classical MPS contains a group of rank-3 tensors $[A^{\delta_{i}}]_{\alpha_{i},\alpha_{i+1}}$ at the $i$-th site. The left and right dimensions ($\alpha_{i}$ and $\alpha_{i+1}$) are auxiliary dimensions representing the correlations between each site. The $N$-qubit state represented by a MPS is
@@
Figure 1: Quantum Matrix Product State model (qMPS). (a) A site of qMPS is a quantum state divided into three entangled groups of qubits, labeled by $l$, $p$ and $r$. Qubits in groups $l$ and $r$ are used as auxiliary dimensions of the qMPS, and the qbits in group $p$ are used ad the physical dimensions of the qMPS. (b) The global state represented by a qMPS is determined by contracting all quantum dimensions, as represented by Eq. 4. The number of qubits contained in the global states is the sum of qubits in physical dimension of each site. (c) The MPO sites can be divided into $m$ groups to match the number of physical qubits in each qMPS site. (d) The local expectation is represented by a rank-6 classical tensor, with each element calculated through Eq. 6.
A qMPS is a multi-component structure that represents a large global state through a group of $m$ distributed smaller-scale quantum states, as shown in Fig. 1. The $i$-th site of the qMPS is a physical quantum state given by
@@
where qubits are partitioned into three entangled groups labeled $l$ (left auxiliary), $p$ (physical) and $r$ (right auxiliary). A qMPS site is naturally a rank-$n_{i}$ quantum tensor Yuan et al. (2021), and the elements of this tensor are the amplitudes of this state. For consistency with a MPS with a bond dimension of $\chi$, we allocate $\lceil\log_{2}(\chi)\rceil$ qubits per auxiliary dimension, and there can be totally $2\lceil\log_{2}(\chi)\rceil+n_{p}$ qubits where $n_{p}$ is the number of physical qubits used in a qMPS site. Generally $n_{p}\geq 1$, and we focus on cases where $n_{p}=1$ to maintain consistency with the classical cases. The data in this site consist of its amplitudes and can be indexed by projecting the state onto the corresponding basis. The space complexity of qMPS naturally scales logarithmically compared to classical MPS.
Contracting two qMPS sites follows the rules Yuan et al. (2021) of
$|\psi_{i,i+1}\rangle=\sum_{\alpha_{i+1}}\langle\alpha_{i+1}|_{r}|\psi_{i}\rangle\otimes\langle\alpha_{i+1}|_{l}|\psi_{i+1}\rangle.\$ (3)
@@
and these qMPS sites form a global $N$-qubit state represented by
$|\Psi\rangle=\sum_{\alpha_{1},\alpha_{2},...,\alpha_{n}}\bigotimes_{i=0}^{m}\langle\alpha_{i}|_{l}\langle\alpha_{i+1}|_{r}|\psi_{i}\rangle,$ (4)
@@
also as shown in Fig. 1(b). The number of qubits in the global state is the sum of qubits in the physical dimension of each qMPS site. The auxiliary dimension, if not bounded, grows exponentially within a classical MPS, while the number of auxiliary qubits grows linearly. Therefore, the qMPS model can be used for strongly entangled systems.
@@
Calculating the global expectation of a qMPS over a global observable can be done locally. In the language of MPS, the observable can first be decomposed into the form of matrix product operator (MPO), and the corresponding physical dimensions are divided into $m$ groups in correspondence to every qMPS site. For any $N$-qubit observable, it can be written as the sum of Pauli products $\hat{O}=\sum_{i=1}^{k}c_{i}(\otimes_{j=1}^{N}P_{ij})$ where $P_{ij}\in\{I,X,Y,Z\}$, and further in a diagonalized MPO form
$\hat{O}$ $=\prod_{i=1}^{N}\hat{O}_{i}$ (5)
$=\begin{bmatrix}P_{11}&\cdots&P_{1k}\end{bmatrix}\begin{bmatrix}P_{21}&&\\ &\ddots&\\ &&P_{2k}\\ \end{bmatrix}\cdots\begin{bmatrix}c_{1}P_{N1}\\ \vdots\\ c_{k}P_{Nk}\end{bmatrix}$
@@
Note that the multiplications between Pauli operators are tensor products. Practically, we only need to record the appearance of each Pauli operator instead of storing every whole MPO site. For simplicity, the coefficients for each term is concentrated in $\hat{O}_{N}$. These $N$ MPO sites can be divided into groups to match the number of physical qubits in each qMPS site, as shown in Fig. 1(c).
The calculation of expectation can be done locally on quantum devices, as shown in Fig. 1(d). Suppose we are now at site $i$ of the qMPS, which is in state $|\psi_{i}\rangle$. The corresponding MPO site is represented by $[\hat{O}^{\sigma_{i},\sigma^{\prime}_{i}}]_{\beta_{i},\beta_{i+1}}$ where $\sigma_{i}$ and $\sigma^{\prime}_{i}$ are the physical dimensions. For a MPO generated following Eq. 5, it satisfies that $[\hat{O}^{\sigma_{i},\sigma^{\prime}_{i}}]_{\beta_{i},\beta_{i+1}}={\bf 0}$ given $\beta_{i}\neq\beta_{i+1}$, and thus is compatible with quantum measurement. The result of the local expectation is therefore a sparse rank-6 classical tensor
@@
$\begin{array}[]{ll}&[V_{i}]_{\alpha_{i},\beta_{i},\alpha^{\prime}_{i}}^{\alpha_{i+1},\beta_{i+1},\alpha^{\prime}_{i+1}}\\ =&\langle\psi_{\alpha^{\prime}_{i},\alpha^{\prime}_{i+1}}^{i}|\hat{O}_{\beta_{i},\beta_{i+1}}|\psi_{\alpha_{i},\alpha_{i+1}}^{i}\rangle\\ =&\langle\psi_{i}|\left(|\alpha_{i}\rangle_{l}\langle\alpha^{\prime}_{i}|_{l}\otimes\hat{O}_{\beta_{i},\beta_{i+1}}\otimes|\alpha_{i+1}\rangle_{r}\langle\alpha^{\prime}_{i+1}|_{r}\right)|\psi_{i}\rangle\end{array}$ (6)
@@
This can be done by measuring $|\psi_{i}\rangle$ by expanding every $|\alpha_{i}\rangle_{l}\langle\alpha^{\prime}_{i}|_{l}$ and $|\alpha_{i+1}\rangle_{r}\langle\alpha^{\prime}_{i+1}|_{r}$ into series of Pauli terms Yuan et al. (2021). Running over $I$, $X$, $Y$ and $Z$ for each connecting qubits and leading to a time complexity of $O(\chi^{4})$ for a single term. The global expectation value can be calculated by classical tensor contraction, and the overall complexity to evaluate the global expectation is $O(km\chi^{4})$, where $k$ is the number of terms in the global observable, and $m$ is the number of qMPS sites.

### v2

@@ Abstract
Solving quantum many-body systems is one of the most significant regimes where quantum computing applies. Currently, as a hardware-friendly computational paradigms, variational algorithms are often used for finding the ground energy of quantum many-body systems. However, running large-scale variational algorithms is challenging, because of the noise as well as the obstacle of barren plateaus. In this work, we propose the quantum version of matrix product state (qMPS), and develop variational quantum algorithms to prepare it in canonical forms, allowing to run the variational MPS method, which is equivalent to the Density Matrix Renormalization Group method, on near term quantum devices. Compared with widely used methods such as variational quantum eigensolver, this method can greatly reduce the number of qubits required, and thus can mitigate the effects of Barren Plateaus while obtain comparable or even better accuracy. Our method holds promise for distributed quantum computing, offering possibilities for fusion of different computing systems.
@@ I Introduction
Solving quantum many-body systems is a fundamental challenge. Classical computers usually have to deal with the exponential growth in size of the system to mimic the quantum effects. As Feynman raised the concept of quantum simulations Feynman1982 , quantum computing begins to show its great power for specific tasks and stimulated the possibility of solving these systems with advantages over the classical computers. Quantum algorithms were proposed for efficient quantum simulations, and are promising for a wide range of applications such as condensed matter physics Anderson1984 ; Sachdev2023 , quantum chemistry Cao2019 ; McArdle2020 , high-energy physics Zohar2016 ; Martinez2016 , materials Monroe2021 ; Maskara2025 , drug discovery Wong2023 ; Santagati2024 , or combinatory optimization tasks Moll2018 ; Harrigan2021 .
@@
As the quantum advantages have been demonstrated on both superconducting and photonic platforms Arute2019 ; Zhu2022 ; Madsen2022 ; Zhong2021 ; Deng2023 ; Acharya2025 ; Gao2025 , quantum computing has developed rapidly and is moving towards the noisy intermediate scale quantum (NISQ) era Preskill2018 . However, taking the practical physical devices into consideration, running large-scale quantum algorithms still meets severe challenges. For example, algorithms such as the variational quantum eigensolver (VQE) Peruzzo2014 ; Kandala2017 , are troubled with ‘Barren Plateaus’ Liu2022 ; Wang2021 , and are more time consuming to converge to an accurate result. To mitigate these challenges, various distributed quantum frameworks have been proposed. For example, deepVQE can handle large-scale systems with weak correlations among subsystems Fujii2022 . Other methods such as circuit cutting Peng2020 ; Tang2022 or hybrid tree tensor networks Yuan2021 reduce the sizes of quantum circuits, but requires global optimizations, which limits the efficiency.
@@
Fortunately, the hybrid tensor network provides a general framework for transplanting the classical tensor network method onto quantum devices. Tensor networks are powerful numerical tools for solving fundamental problems in quantum many-body systems. Recently, tensor network methods are applied in quantum information, and become one of the most efficient method for simulating quantum circuits Markov2018 ; Villalonga2019 ; Villalonga2020 ; GuoLiu2019 ; Liu2021 ; Pan2022 . As one of the most widely used tensor network structures, Matrix Product State (MPS) supports important applications of finding the ground energy of a strongly correlated 1-D many-body system Ostlund1995 ; Hastings2007 ; Liu2020 . This tensor network structure inspires a quantum circuit ansatz for data encoding or further processing Ran2020 ; Dilip2022 ; Jaderberg2025 . We focus on a crucial feature of MPS: renormalizing the many-body system and breaking it down into several subtasks executable on quantum devices.
@@
In this work, we propose the idea of quantum Matrix Product State (qMPS), and develop the variational quantum MPS (vqMPS) method for quantum simulation, which can be viewed as the quantum version of the Density Matrix Renormalization Group (DMRG) method White1992 ; Schollwock2011 for near-term quantum devices. This method makes it possible to run large-scale quantum simulations with local optimizations on small-scale subsystems, and can reach comparable or even better accuracy over VQE with limited qubits. This can greatly enhance the efficiency for both physical demonstration and numerical simulation. Also, the local optimizations can be distributed on different systems, providing a framework for the fusion of different physical platforms, or further quantum-classical hybrid computing.
@@ II The quantum MPS
A classical MPS contains a group of rank-3 tensors $A^{\delta_{i}}]_{\alpha_{i},\alpha_{i+1}}$ at the $i$-th site. The left and right dimensions ($\alpha_{i}$ and $\alpha_{i+1}$) are auxiliary dimensions representing the correlations between each site. The $N$-qubit state represented by a MPS is
@@
Figure 1: Quantum Matrix Product State model (qMPS). (a) A diagram of implementing a qMPS site on a superconducting processor. A site of qMPS is a quantum state divided into three entangled groups of qubits, labeled by $l$, $p$ and $r$. Qubits in groups $l$ and $r$ are used as auxiliary dimensions of the qMPS, and the qubits in group $p$ are used ad the physical dimensions of the qMPS. (b) The global state represented by a qMPS is determined by contracting all quantum dimensions, as represented by Eq. 4. The number of qubits contained in the global states is the sum of qubits in physical dimension of each site. (c) The MPO sites can be divided into $m$ groups to match the number of physical qubits in each qMPS site. (d) The local expectation is represented by a rank-6 classical tensor, with each element calculated through Eq. 7.
A qMPS is a multi-component structure that represents a large global state through a group of $m$ distributed smaller-scale quantum states. The $i$-th site of the qMPS is a physical quantum state containing $n_{i}$ qubits, which is given by
@@
where qubits are partitioned into three entangled groups labeled $l$ (left auxiliary), $p$ (physical) and $r$ (right auxiliary). Fig. 1(a) is a diagram of implementing a qMPS site on a superconducting processor. A qMPS site is naturally a rank-$n_{i}$ quantum tensor Yuan2021 , and the elements of this tensor are the amplitudes of this state. For consistency with a MPS with a bond dimension of $\chi$, we allocate $n_{l}=n_{r}=\lceil\log_{2}(\chi)\rceil$ qubits per auxiliary dimension, and there can be totally $2\lceil\log_{2}(\chi)\rceil+n_{p}$ qubits where $n_{p}$ is the number of physical qubits used in a qMPS site. Generally $n_{p}\geq 1$, and we focus on cases where $n_{p}=1$ to maintain consistency with the classical cases. The data in this site consist of its amplitudes and can be indexed by projecting the state onto the corresponding basis. The space complexity of qMPS naturally scales logarithmically compared to classical MPS.
Contracting two qMPS sites follows the rules Yuan2021 of
$|\psi_{i,i+1}\rangle=$ $\sum_{\alpha_{i+1}}\langle\alpha_{i+1}|_{r}\cdot|\psi_{i}\rangle\otimes\langle\alpha_{i+1}|_{l}\cdot|\psi_{i+1}\rangle$ (3)
$=$ $\sum_{\alpha_{i+1}}\left(\sum_{\alpha_{i},\alpha_{i+2}}c^{i}_{\alpha_{i},\alpha_{i+1}}c^{i+1}_{\alpha_{i+1},\alpha_{i+2}}|\alpha_{i}\rangle_{l}|\psi_{\alpha_{i},\alpha_{i+1}}^{i}\rangle_{p}|\psi_{\alpha_{i},\alpha_{i+1}}^{i}\rangle_{p}|\alpha_{i+1}\rangle_{r}\right),$
@@
These qMPS sites collectively form a global $N$-qubit state, as represented by equation:
$|\Psi\rangle=\sum_{\alpha_{1},\alpha_{2},...,\alpha_{n}}\bigotimes_{i=0}^{m}(\langle\alpha_{i}|_{l}\langle\alpha_{i+1}|_{r})\cdot|\psi_{i}\rangle,$ (4)
@@
also as shown in Fig. 1(b). The number of qubits in the global state is the sum of the number of qubits representing the physical dimension of each qMPS site. The auxiliary dimension, if not bounded, grows exponentially within a classical MPS, while the number of auxiliary qubits grows linearly. Therefore, the qMPS model can be used for strongly entangled systems.
@@
Calculating the global expectation of a qMPS for a global observable can be done through local measurements. In the language of MPS, the observable can first be decomposed into a matrix product operator (MPO), and the corresponding physical dimensions are divided into $m$ groups in correspondence to every qMPS site. For any $N$-qubit observable, it can be written as the sum of Pauli products $\hat{O}=\sum_{i=1}^{K}c_{i}(\otimes_{j=1}^{N}P_{ij})$ where $P_{ij}\in\{I,\sigma_{x},\sigma_{y},\sigma_{z}\}$, and further in a simplified and diagonalized MPO form
$\hat{O}=[\hat{O}_{1}]\times[\hat{O}_{2}]\times\cdots\times[\hat{O}_{N}]=\begin{bmatrix}P_{11}&\cdots&P_{1K}\end{bmatrix}\times\begin{bmatrix}P_{21}&&\\ &\ddots&\\ &&P_{2K}\\ \end{bmatrix}\times\cdots\times\begin{bmatrix}c_{1}P_{N1}\\ \vdots\\ c_{K}P_{NK}\end{bmatrix}$. (5)
@@
Note that the multiplications between Pauli operators are tensor products. In practice, it is sufficient to record the appearance of each Pauli operator instead of storing a huge amount of MPO data. For simplicity, the coefficients for each term are concentrated in $\hat{O}_{N}$. These $N$ MPO sites can be divided into groups to match the number of physical qubits in each qMPS site, as shown in Fig. 1(c), while in this work, we focus on the cases where $m=K$, i.e., the MPO sites are separated per physical qubit of the qMPS sites.
The calculation of expectation can be done site-by-site and locally on a smaller quantum device, as shown in Fig. 1(d). Suppose we are now at site $i$ of the qMPS, which is in state $|\psi_{i}\rangle$. The corresponding MPO site is represented by $\hat{O}^{\sigma_{i},\sigma^{\prime}_{i}}]_{\beta_{i},\beta_{i+1}}$ where $\sigma_{i}$ and $\sigma^{\prime}_{i}$ are the physical dimensions, while $\beta_{i}$ and $\beta_{i+1}$ are assistant dimensions linking to the neighbouring sites. As shown in Fig. 1(d), the MPO site only covers the physical dimensions of $|\psi_{i}\rangle$. The auxiliary qubits act as free indices, and can be transferred into classical indices by projecting the auxiliary qubits onto the computational basis, obtaining a list of (unnormalized) quantum states as
$|\psi_{i}(j,k)\rangle$ $=\langle j|_{l}\langle k|_{r}\cdot|\psi_{i}\rangle$ (6)
$=\langle j|_{l}\langle k|_{r}\left(\sum_{\alpha_{i},\alpha_{i+1}}c^{i}_{\alpha_{i},\alpha_{i+1}}|\alpha_{i}\rangle_{l}|\psi_{\alpha_{i},\alpha_{i+1}}^{i}\rangle_{p}|\alpha_{i+1}\rangle_{r}\right)$
$=c_{j,k}^{i}|\psi_{j,k}^{i}\rangle.$
For an MPO generated following Eq. 5, it satisfies that $\hat{O}^{\sigma_{i},\sigma^{\prime}_{i}}]_{\beta_{i},\beta_{i+1}}=\delta_{\beta_{i},\beta_{i+1}}P_{i,\beta_{i}}$, which indicates that $\hat{O}^{\sigma_{i},\sigma^{\prime}_{i}}]_{\beta_{i},\beta_{i+1}}={\bf 0}$ given $\beta_{i}\neq\beta_{i+1}$ and we can only iterate over the pauli operators appearing in its diagonal. Therefore, it also provides two free indices, and as a result, the local expectation is a rank-6 tensor (or a rank-3 tensor if the site is at the end of the chain), as shown in eq. 7.
@@
$\begin{array}[]{ll}[V_{i}]_{\gamma_{i},\beta_{i},\gamma^{\prime}_{i}}^{\gamma_{i+1},\beta_{i+1},\gamma^{\prime}_{i+1}}&=\langle\psi_{i}(\gamma^{\prime}_{i},\gamma^{\prime}_{i+1})|\hat{O}_{\beta_{i},\beta_{i+1}}|\psi_{i}(\gamma_{i},\gamma_{i+1})\rangle\\ &=\left(\langle\psi_{i}|\cdot|\gamma^{\prime}_{i}\rangle_{l}|\gamma^{\prime}_{i+1}\rangle_{r}\right)\hat{O}_{\beta_{i},\beta_{i+1}}\left(\langle\gamma_{i}|_{l}\langle\gamma_{i+1}|_{r}\cdot|\psi_{i}\rangle\right)\\ &=\langle\psi_{i}|\left(|\gamma^{\prime}_{i}\rangle_{l}\langle\gamma_{i}|_{l}\otimes\hat{O}_{\beta_{i},\beta_{i+1}}\otimes|\gamma^{\prime}_{i+1}\rangle_{r}\langle\gamma_{i+1}|_{r}\right)|\psi_{i}\rangle.\end{array}$ (7)
@@
In practice, $\hat{O}_{\beta_{i},\beta_{i+1}}$ is a legal observable, and each $|\gamma_{i}\rangle_{l}\langle\gamma^{\prime}_{i}|_{l}$ or $|\gamma_{i+1}\rangle_{r}\langle\gamma^{\prime}_{i+1}|_{r}$ can be expanded into series of Pauli terms. Then, the local expectation can be measured experimentally. Running over $I$, $\sigma_{x}$, $\sigma_{y}$ and $\sigma_{z}$ for each auxiliary qubit leads to a time complexity of $O(\chi^{4})$ for a single term. The global expectation value can be calculated by classical tensor contraction, and the overall complexity to evaluate the global expectation is $O(Km\chi^{4})$, where $K$ is the number of terms in the global observable, and $m$ is the number of qMPS sites. Once all local expectations are measured, the global expectation can be calculated through classical tensor contraction
$V=\sum_{\begin{subarray}{c}\gamma_{1},\gamma_{2},...,\gamma_{m+1}\\ \beta_{1},\beta_{2},...,\beta_{m+1}\\ \gamma^{\prime}_{1},\gamma^{\prime}_{2},...,\gamma^{\prime}_{m+1}\end{subarray}}[V_{1}]_{\gamma_{1},\beta_{1},\gamma^{\prime}_{1}}^{\gamma_{2},\beta_{2},\gamma^{\prime}_{2}}[V_{2}]_{\gamma_{2},\beta_{2},\gamma^{\prime}_{2}}^{\gamma_{3},\beta_{3},\gamma^{\prime}_{3}}\cdots[V_{m}]_{\gamma_{m},\beta_{m},\gamma^{\prime}_{m}}^{\gamma_{m+1},\beta_{m+1},\gamma^{\prime}_{m+1}}.$ (8)